OpenStudy (anonymous):

Which calculation can be used to find the value of p in the equation p3 = 8? p equals cube root of 8 p equals 3 multiplied by the square root of 8 p equals 2 multiplied by square root of 8 p equals square root of 8

OpenStudy (anonymous):

\[p3=\sqrt{8}\] is that how its supposed to look?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

or \[\frac{ p }{ 3 }=\sqrt{8}\]

OpenStudy (anonymous):

p3 = 8?

OpenStudy (anonymous):

its between a and d

OpenStudy (anonymous):

but i think a more